OpenStudy (anonymous):

Graph the linear inequality. y<1/5x+6

pooja195 (pooja195):

Lets determine weather our line will be dashed or dotted any ideas?

pooja195 (pooja195):

*whether

OpenStudy (anonymous):

i think solid line is < or > correct me if I am wrong lol

pooja195 (pooja195):

Incorrect it would be a dashed line because a solid line is used for \[\le ~or \ge \]

OpenStudy (anonymous):

okay so how do we do this

pooja195 (pooja195):

Ok When our starting point is since the y intercept is 6

pooja195 (pooja195):

The site is lagging

pooja195 (pooja195):



pooja195 (pooja195):

Starting point is 6

pooja195 (pooja195):



OpenStudy (anonymous):

is that the answer I am confused, sorry

pooja195 (pooja195):

Well its really hard to show graphing problems on here when the site lags so yes this is the graph of the inequality It starts at 6 all values less than 6 are shaded.

OpenStudy (anonymous):

okay, i apoligize

pooja195 (pooja195):

Its alright :) its not your fault but the graph is the answer

pooja195 (pooja195):

Do you understand though?

OpenStudy (anonymous):

yes I do now thanks

pooja195 (pooja195):

np =)

TheSmartOne (thesmartone):

Here's the actual way you would do it, if you were doing it in a class: You first graph the line y = (1/5)x + 6 And before you graph it, you have to determine whether the graph is dotted or full. In our case it is dotted. Dotted/dashed same thing :P Next you plug in a random point into the equation y < (1/5)x + 6 to see which side you should shade.

TheSmartOne (thesmartone):

If you need me to go more in-depth, just ask :)